# Diffusion models for multivariate subsurface generation and efficient probabilistic inversion

### v1

@@ 4 Discussion
The trained DM effectively samples from complex bivariate geological priors. To evaluate performance, we mainly focused on first- and second- order statistics (higher-order aspects is considered in Appendix A). The generated images match the TI features with a fidelity that, to our knowledge, is unprecedented for multivariate geological modeling with deep generative models. For the considered metrics, the DM outperformed the VAE and GAN implementations by Miele et al. (2024) (e.g., Fig. 2 and Table 1), except for a general overestimation of sand volume fraction that was slightly less pronounced in the VAE results (Table 1 and Figure S1 in Appendix A). A practical advantage of DMs for multivariate modeling is the possibility to associate individual channels of the underlying U-Net to specific subsurface property types without the need to modify the network architecture. In contrast, VAE and GAN modeling required independent convolutional layers for each property type. This suggests that pre-designed DM architectures require none or limited adjustments for modeling different geological scenarios.
@@
Using the proposed CDPS, pre-trained DMs can be flexibly modified to achieve noise-robust posterior sampling in both linear and non-linear inversion settings. The framework can be used for a wide range of inversion tasks without retraining, and deal with the simultaneous use of multiple conditioning data types (Appendix B). The method includes modifications to the DPS by Chung et al. (2023), mainly an improved likelihood score approximation. By accounting for the error in $\hat{\textbf{x}}_{0}^{(t)}$ on the simulated conditioning data (Eq. 19), the likelihood scores become more accurate and scaled to diffusion noise (Fig. 5). With this modification, we allow a much more stable generative process compared to DPS, balancing the two scores contributions for posterior sampling (Eq. 10) automatically (i.e., without arbitrary weighting), and making the likelihood score more independent of the data noise. However, a small noise-dependency remains for the nonlinear case with CDPS (Fig. 5b). This might be due to the overly simple assumption of homoscedastic noise in the $\hat{\textbf{x}}_{0}^{(t)}$ estimates (Fig. 4), which may not hold if the trained network presents any form of localized bias (due to wrong training procedure, network’s architecture and/or biased TI distributions). Improved error estimates may be needed in future applications, considering locally dependent variances or using approximations of Eq. 15.
@@
Our proposed CDPS algorithm (Algorithm 1) enables efficient and more accurate local conditioning than DPS (see, e.g., Fig. 6a and b). We did not manage to properly benchmark the results against MCMC with pCN proposals (Fig. 6c). Indeed, even in the linear inversion case there is a relatively poor mixing and limited convergence of pCN; we did not investigate more advanced MCMC approaches such as parallel tempering formulations (Xu et al., 2020). However, there is an overall good agreement between CDPS and pCN samples at the well locations (Fig. 7 and Table 3). This suggests that CDPS appropriately approximate the posterior. Compared with DPS (Fig. 7b), CDPS (Fig. 7a) has wider uncertainty bounds of magnitudes similar to those obtained by MCMC. We postulate that the underestimation by DPS is primarily a consequence of the likelihood score magnitudes being overestimated (Fig. 5a) when using the approximation in Eq. 12.
@@
For the non-linear inversion case, the CDPS performed well for all considered noise levels while DPS failed when considering low data noise levels (e.g., Fig. 6 and Table 5). The DPS used herein is an implementation of the theory by Chung et al. (2023) using the score-based EDM framework. We also considered arbitrarily-weighted Gaussian likelihood functions or the RMSE for the log-score approximation (as in the original DDIM implementation); this led to no convergence and unstable behavior in all the considered cases. We reproduce similar results using CDPS with DDIM (Appendix C) highlighting that the method also works well for this formulation.
@@
Even if the CDPS method performs very well in practice, it is an approximate method because the logarithmic score estimate (Eq. 19) relies on an approximation of the intractable likelihood (Eq. 11), the prior score estimates provided by the neural network will inevitably have some errors and the number of diffusion time steps are finite. One computational challenge of the CDPS method is the need to estimate the Jacobian of the forward operator at each time step. When this calculation is very expensive, we anticipate that the CDPS would still work well even if the Jacobian is only updated occasionally during the generative process.
@@ 5 Conclusions
We investigate diffusion models (DM) for sampling multivariate geological priors and propose a modified method (CDPS) for approximate probabilistic Bayesian inversion. itsperformance is evaluated on a bivariate distribution of lenticular sand deposits and corresponding acoustic impedances ($I_{P}$). The unconditional DM simulations effectively captures first- and second-order statistics, as well as higher-order facies structures represented in training images (TI), consistently outperforming benchmark generative models based on VAEs and GANs. In contrast to these methods, the DM does not require architectural modifications for modeling bivariate geological properties.
The inversion method leverages the iterative nature of diffusion modeling; posterior samples are obtained by considering, at each iteration of the denoising process, a likelihood score approximation that modify the update direction during the generative process. There is no need to train the DM when considering conditioning data and the inversion is performed without any form of data dimensionality reduction. We demonstrate that the method enables efficient Bayesian posterior sampling for both linear (borehole data) and nonlinear (fullstack seismic) data scenarios, or combinations thereof. Our implementation makes corrections to the likelihood approximation of the original Diffusion Posterior Sampling (DPS) method (DPS; Chung et al., 2023). This leads to stabilized inversion across noise levels, as well as more accurate posterior estimates indicated by lower bias and better exploration of the posterior pdf. For the considered test examples, CDPS also shows significant advantages over an MCMC formulation in terms of convergence and vastly decreased computing times, particularly for nonlinear and high-dimensional data settings.
This work highlights the strengths of diffusion-based generative models for complex subsurface modeling and inference tasks. The training is much more robust than for GANs and the data generation quality is beyond the capabilities of GANs and VAEs. The CDPS approach enables flexible and scalable probabilistic inversion with minimal architectural tuning, resulting in an approximate Bayesian inversion method with broad applications across complex geological scenarios and data types.
@@ A.1 Geostatistical metrics
Figure 9: Volume fraction of sands in the generated realizations, compared to TI (Prior), for the fully trained models (a), and for early training stages at epoch 100 (b).
@@ A.2 Training efficiency
Table 7: Metrics for unconditional prior modeling using the trained DM and benchmarking GAN and VAE at epoch 100. p(sand) and Avg. $I_{P}$ refer to the average models obtained from the realizations ensembles. $KL$: Kullback–Leibler divergence; Avg.: average; St. dev.: standard deviation.
@@
Figure 12: Summary of unconditional modeling performances of DM (a, b), GAN (c, d), and VAE (e, f) at training epoch 100. Subplots (a), (c), (e): realizations and average distribution of facies and $I_{P}$ for DM (a), GAN (c) and VAE (e); subplots (b), (d), (f): joint and $I_{P}$ distributions for DM (b), GAN (d) and VAE (f).
@@ Appendix B Conditioning on multiple data types
Figure 13: Target facies and $I_{P}$ for three case studies (a, b, c), sampled posterior distribution shown as mean and standard deviation (Std. dev.) of realizations’ ensembles, and corresponding average seismic data misfit.
Figure 14: Predicted $I_{P}$ and corresponding seismic data vertical traces at well location (x = 50 m) and at x = 25 m, compared with conditioning data $\bf{d_{obs}}$ and unknown true $I_{P}$ (Target).

### v2

@@ 4 Discussion
The trained DM effectively samples from complex bivariate geological priors. To evaluate performance, we mainly focused on first- and second-order statistics (higher-order features are considered in the Appendix A). The generated images match the TI features with a fidelity that, to our knowledge, is unprecedented for multivariate geological modeling with deep generative models. For the considered metrics, the DM outperformed the VAE and GAN implementations by Miele et al. (2024) (e.g., Fig. 2 and Table 1), except for a general overestimation of sand volume fraction that was slightly less pronounced in the VAE results (Table 1 and Figure S1 in Appendix A). A practical advantage of DMs for multivariate modeling is the possibility to associate individual channels of the underlying U-Net to specific subsurface property types without the need to modify the network architecture. In contrast, VAE and GAN modeling required independent convolutional layers for each property type. This suggests that pre-designed DM architectures require none or only limited adjustments for modeling different geological scenarios.
@@
Using the proposed CDPS, pre-trained DMs can be flexibly modified to achieve noise-robust posterior sampling in both linear and non-linear inversion settings. The framework can be used for a wide range of inversion tasks without retraining, and it can deal with the simultaneous use of multiple conditioning data types (Appendix B). The method includes modifications to the DPS by Chung et al. (2023), mainly in terms of an an improved likelihood score approximation. By accounting for the error in $\hat{\textbf{x}}_{0}^{(t)}$ on the simulated conditioning data (Eq. 19), the likelihood scores become more accurate and scaled to diffusion noise (Fig. 5). With this modification, we obtain a much more stable generative process compared to DPS, balancing the two score contributions for posterior sampling (Eq. 10) automatically (i.e., without arbitrary weighting), and making the likelihood score more independent of the data noise. However, a small noise-dependency remains for the nonlinear case with CDPS (Fig. 5b). This might be due to the overly simple assumption of homoscedastic noise in the $\hat{\textbf{x}}_{0}^{(t)}$ estimates (Fig. 4), which may not hold if the trained network presents any form of localized bias (due to an incorrect training procedure, network’s architecture and/or biased TI distributions). Improved error estimates may be needed in future applications, considering locally dependent variances or using approximations of Eq. 15.
@@
Our proposed CDPS algorithm (Algorithm 1) enables efficient and more accurate local conditioning than DPS (see, e.g., Fig. 6a and b). We did not manage to properly benchmark the results against MCMC with pCN proposals (Fig. 6c). Indeed, even in the linear inversion case there is a relatively poor mixing and limited convergence of pCN MCMC; we did not investigate more advanced MCMC approaches such as parallel tempering formulations (Xu et al., 2020). However, there is an overall good agreement between CDPS and pCN samples at the well locations (Fig. 7 and Table 3). This suggests that CDPS appropriately approximate the posterior. Compared with DPS (Fig. 7b), CDPS (Fig. 7a) has wider uncertainty bounds with magnitudes similar to those obtained by MCMC. We postulate that the underestimation by DPS is primarily a consequence of the likelihood score magnitudes being overestimated (Fig. 5a) when using Eq. 12. For the non-linear inversion case, the CDPS performed well for all considered noise levels while DPS failed when considering low data noise levels (e.g., Fig. 5 and Table 5). Similarly to the linear case, we attribute the unstable DPS denoising trends (Fig. 5) to overestimated likelihood scores, with the overestimation increasing with decreasing data noise.
@@
The DPS used herein is an implementation of the theory by Chung et al. (2023) using the score-based EDM framework. We also considered arbitrarily-weighted Gaussian likelihood functions or the RMSE for the log-score approximation (as in the original DDIM implementation); this led to no convergence and unstable behavior in all the considered cases. We reproduce similar results using CDPS with DDIM (Appendix C) highlighting that the method also works well for this formulation.
@@
Even if the CDPS performs well in practice, it is still an approximate method because the logarithmic score estimate (Eq. 19) relies on an approximation of the intractable likelihood (Eq. 11), the prior score estimates provided by the neural network will inevitably have some errors and the number of diffusion time steps are finite. One computational challenge of the CDPS method is the need to estimate the Jacobian of the forward operator at each time step. When this calculation is expensive, we anticipate that the CDPS approach would still work well in practice even if the Jacobian is only updated occasionally during the generative process.
@@ 5 Conclusions
We investigate diffusion models (DM) for sampling multivariate geological priors and propose a modified conditional sampling method (CDPS) for approximate probabilistic Bayesian inversion. The method leverages the iterative nature of diffusion modeling, using an additional likelihood score approximation that modifies the generative process, allowing us to use any pre-traind, unconditional DM. The modeling performances are evaluated on a bivariate distribution of lenticular sand deposits and corresponding acoustic impedances ($I_{P}$). The unconditional DM simulations effectively captures first- and second-order statistics, as well as higher-order facies structures represented in training images (TI), consistently outperforming benchmark generative models based on VAEs and GANs. Bayesian posterior sampling is demonstrated for both linear (borehole data) and nonlinear (fullstack seismic) data scenarios, or combinations thereof. Compared to the original Diffusion Posterior Sampling (DPS) method (DPS; Chung et al., 2023), our CDPS implementation provides stabilized inversion across noise levels, as well as more accurate posterior estimates, indicated by lower bias and enhancedr exploration of the posterior pdf. The method also shows significant advantages over an MCMC formulation in terms of convergence, leading to vastly decreased computing times. This work highlights the strengths of diffusion-based generative models for complex subsurface modeling and inference tasks. The training is much more robust than for GANs and the data generation quality is beyond the capabilities of GANs and VAEs. The CDPS approach enables flexible and scalable probabilistic inversion with minimal architectural tuning, resulting in an approximate Bayesian inversion method with broad applications across complex geological scenarios and data types.
@@ A.1 Geostatistical metrics
Figure 9: Volume fraction of sands in the generated realizations, compared to TI (Prior), for (a) the fully trained models, and (b) for early training stages at epoch 100.
@@ A.2 Training efficiency
Table 7: Metrics for unconditional prior modeling using the trained DM and benchmarking GAN and VAE at epoch 100. p(sand) and Avg. $I_{P}$ refer to the average models obtained from the realizations ensembles. KLKL: Kullback–Leibler divergence; Avg.: average; St. dev.: standard deviation.
@@
Figure 12: Summary of unconditional modeling performances of (a, b) DM, (c, d) GAN, and (e, f) VAE at training epoch 100. Subplots (a), (c), (e): realizations and average distribution of facies and $I_{P}$ for (a) DM, (c) GAN and (e) VAE; subplots (b), (d), (f): joint and $I_{P}$ distributions for (b) DM, (d) GAN and (f) VAE.
@@ Appendix B Conditioning on multiple data types
Figure 13: Target and posterior facies and $I_{P}$ for the three inversion case studies (a, b, c) using both well logs and seismic data as conditioning data. The sampled posterior distribution is shown as mean and standard deviation (Std. dev.) of realizations’ ensembles, and corresponding average seismic data misfit.
Figure 14: Predicted $I_{P}$ and corresponding seismic data along vertical traces located at the well location (x = 50 m) and at x = 25 m, for the three inversion case studies (a, b, c). The predictions are compared with conditioning data $\bf{d_{obs}}$ and unknown true $I_{P}$ (Target).